OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = -1.

OpenStudy (anonymous):

A) f(x) = −one twelfth (x − 5)2 + 2 B) f(x) = one twelfth (x − 5)2 + 2 C) f(x) = −one twelfth (x + 5)2 + 2 D) f(x) = one twelfth (x + 5)2 + 2

OpenStudy (mathmale):

Do you think this parabola is a vertical or a horizontal one? How would you know? Note that the vertex of a parabola is located exactly half-way between the focus and the directrix. Please find the coordinates of the focus.

OpenStudy (mathmale):

I'd suggest that you draw a graph, to help you visualize this situation.